### v3

# Bipartite Fluctuations and Charge Fractionalization in Quantum Wires

Abstract
We introduce a quantum information method for measuring fractional charges in ballistic quantum wires generalizing bipartite fluctuations to the chiral quasiparticles in Luttinger liquids, i.e. adding charge and current fluctuations in a region of the wire. Bipartite fluctuations at equilibrium are characterized through a logarithmic scaling with distance encoding the entangled nature of these fractional charges in one-dimensional (1D) fluids. This approach clarifies the physical meaning of the dephasing factor of electronic interferences in a ballistic ring geometry at zero temperature, as a result of charge fractionalization. We formulate an analogy towards ground-state energetics. We show how bipartite current fluctuations represent a useful tool to locate quantum phase transitions associated to Mott physics. We address a spin chain equivalence and verify the fractional charges through an algorithm such as Density Matrix Renormalization Group (DMRG). Adding a potential difference between the two sides (parties) of the wire, bipartite fluctuations can detect a bound state localized at the interface through the Jackiw-Rebbi model coexisting with fractional charges.
preprint: APS/123-QED
Introduction.— In 1D quantum systems, electronic interactions give rise to unique and paradigmatic phenomena [1, 2]. The ground-state wavefunction of a Luttinger liquid has a Jastrow form similar to the Laughlin wavefunction [3, 4]. Low-dimensional ballistic quantum wires then are characterized by charge fractionalization similar to the fractional quantum Hall physics, but chiral quasiparticles can move in both directions, left or right [3, 5]. It is yet delicate to measure those fractional charges. The conductance of a quantum wire reaches $\frac{e^{2}}{h}$ as for free electrons, assuming spin-polarized electrons, which can be understood as a Fabry-Perot resonance of fractional wavepackets [5, 6, 7]. Signatures of charge fractionalization can be found through momentum-resolved tunneling when injecting an electron with fixed chirality in the middle of a wire [8, 9]; this measure reveals the difference or Asymmetry between the (fractional) charges moving in different directions which is proportional to the Luttinger parameter. Interferometry measures in a ballistic wire allows us to access informations on dephasing effects from interactions, e.g. at finite temperature the dephasing time reveals the associated electron lifetime in a Luttinger liquid [10, 11]. A dephasing time proportional to the inverse of temperature was reported in ballistic rings of GaAs [12], but both electron-electron interactions [10] and dephasing from a surrounding metallic gate [13] can explain the signal. In chiral edge states of fractional quantum Hall systems various tools are developed and proposed to measure the fractional charges with great accuracy [14] such as current noise at zero frequency [15, 16] and Ramsey interferometry in time [17].
To motivate further the present work, we observe that direct measures of many-body entangled systems, e.g. through bipartite entanglement entropy [18, 19], are also showing a huge interest in the community building a bridge with quantum information theory. For Luttinger liquids, this entropy shows a logarithmic behavior with distance [20]. The 1D Fermi gas and the Luttinger liquid reveal the same entanglement entropy. In this work, our goal is to clarify how we can reveal fractional charges of 1D quantum wires from such equilibrium methods in a quantum information perspective. From this angle, it is useful to take the point of view of bipartite charge fluctuations [21] that we introduced as a possible marker of the many-body entangled nature of ground states e.g. in Luttinger liquids taking place between two macroscopic regions of the sample; this gives rise to a logarithmic behavior as a function of length or as a function of time similar to the entanglement entropy. Bipartite charge fluctuations mean that we study the variance (noise) associated to the charge or number of electrons in one of the two regions of the wire. This probe reveals the Luttinger parameter $K$ as a prefactor which has the significance of the compressibility from an analogy to statistical physics [21]. To reveal precisely the fractional charges $f_{R}=\frac{1+K}{2}$ and $f_{L}=\frac{1-K}{2}$ in units of the electron charge [3, 5, 8, 9, 10], in this work, we introduce the bipartite chiral charge fluctuations i.e. bipartite fluctuations of charge associated to the true chiral quasiparticles which move to the left and to the right. This will be obtained when adding bipartite charge and current fluctuations: Eqs. (7), (8) and (9) will reveal the fractional charges i.e. $f_{L}^{2}+f_{R}^{2}$ as a prefactor of the logarithmic scaling. From Landauer’s point of view, noise here is a quantum signal i.e. it precisely measures the entangled nature of chiral quasiparticles. We clarify the link with the dephasing of mesoscopic interferences at zero temperature in Eq. (12). Bipartite charge fluctuations of chiral quasiparticles in the fractional quantum Hall effect are also a useful tool to probe edge states [22, 23].
Luttinger liquid and fractional charges.— The Luttinger liquid takes the form of a simple quadratic string model through bosonization [1, 2]
$H=\frac{u}{2\pi}\int_{0}^{L}dx\left(\frac{1}{K}(\partial_{x}\phi)^{2}+K(\partial_{x}\theta)^{2}\right),$ (1)
where the current density takes the form $j=\frac{uK}{\pi}\partial_{x}\theta$ and the charge density reads $\rho=\rho_{0}-\frac{1}{\pi}\partial_{x}\phi$ with $\rho_{0}$ the mean density of electrons. Here, $0<K<1$ is the Luttinger parameter and $u$ the sound velocity. For free electrons $K=1$ and $u$ is equal to the Fermi velocity $v_{F}$. We can then introduce the total charge $Q$ (normalized to $e$) and total current $J$ such that $Q=\int_{0}^{L}dx\rho(x,t)=\rho_{0}L-\frac{1}{\pi}(\phi(L)-\phi(0))$ and $J=\int_{0}^{L}dxj(x,t)=\frac{uK}{\pi}(\theta(L)-\theta(0))=uK\tilde{J}$ with $uK=v_{F}$. The wire has a length $L$ and for simplicity we study spin-polarized electrons to probe the charge dynamics. On the lattice, we have the correspondences $\sum_{i=0}^{L}c^{\dagger}_{i}c_{i}=\rho_{0}L-\frac{1}{\pi}(\phi(L)-\phi(0))$ and $\sum_{i=0}^{L}\left(c^{\dagger}_{i}c_{i-1}-c^{\dagger}_{i-1}c_{i}\right)=-\frac{2i}{\pi}(\theta(L)-\theta(0))$. The Luttinger Hamiltonian corresponds to interacting fermions hopping on a lattice through adjacent sites [1, 2].
These definitions are particularly meaningful to introduce the chiral quasiparticles which propagate at the sound velocity $u$ in the fluid. In this way, we have the identifications for the charge and current densities $\rho=n_{L}+n_{R}$ and $j=u(n_{R}-n_{L})$ in terms of the chiral densities at position $x$ moving to the right $n_{R}$ and to the left $n_{L}$ respectively. We can similarly introduce $Q_{L}=\int_{0}^{L}n_{L}(x)dx$ and $Q_{R}=\int_{0}^{L}n_{R}(x)dx$. We obtain the usual identities $Q=Q_{R}+Q_{L}$ and $J=u(Q_{R}-Q_{L})$, the charge of an electron is unity. The last equality is equivalent to $\tilde{J}=\frac{J}{uK}$ within our definitions i.e. $K\tilde{J}=Q_{R}-Q_{L}$. Therefore, suppose we have a right-moving electron injected from a reservoir with the Fermi velocity; this will give rise to the identities $Q=+1$ and $J=+v_{F}=+uK$ or $\tilde{J}=+1$ and therefore to chiral fractional charges moving to the right and to the left $Q_{R}=\frac{1+K}{2}=f_{R}$ and $Q_{L}=\frac{1-K}{2}=f_{L}$. Injecting an electron with the Fermi velocity means that the reservoir can be seen as a 1D Fermi gas with a Luttinger parameter $K=1$ [5, 9]. Similarly, we can introduce a left-moving electron from a reservoir corresponding then to $J=-v_{F}=-uK$ i.e. to $\tilde{J}=-1$ corresponding to invert the formulas for $f_{L}$ and $f_{R}$ i.e. $f_{L}=\frac{1+K}{2}$ and $f_{R}=\frac{1-K}{2}$. Momentum-resolved tunneling reveals the difference of fractional charges $f_{R}-f_{L}=\pm K$ [8, 9]. Bipartite fluctuations associated to the chiral quasiparticles will reveal the prefactor $\frac{1}{K}(f_{L}^{2}+f_{R}^{2})=\frac{1}{2}(K+K^{-1})$ which is identical for electrons propagating to the left and to the right.
The Hamiltonian can be equivalently written as
$H=\frac{uK}{4\pi}\int_{0}^{L}dx\left((\partial_{x}\theta_{R}(x))^{2}+(\partial_{x}\theta_{L}(x))^{2}\right),$ (2)
where $\theta_{R,L}=\theta\mp\frac{\phi}{K}$ [3, 10]. The associate charge variables read $\phi_{R,L}=\frac{1}{2}(-\phi\pm K\theta)$ such that $\phi=(\phi_{R}+\phi_{L})$ and $\theta=(\phi_{R}-\phi_{L})/K$. We remind the commutation relations $\phi(x),\nabla\theta(y)]=i\pi\delta(y-x)$. Integrating the chiral charge densities $\partial_{x}\phi_{R,L}$, then this reveals
$\int_{0}^{L}\partial_{x}\phi_{R,L}dx=\frac{Q\pm K\tilde{J}}{2}=Q_{R,L}.$ (3)
It is also possible to introduce quasiparticle operators for the fractional charges associated to the electron operator [3] ${\cal L}_{R,L}^{f_{R},f_{L}}(x)=\exp(-if_{R,L}\theta_{R,L}(x))$ which gives a physical interpretation to the electron Green’s function and to the electron lifetime [10].
To study bipartite fluctuations we introduce the charge and current operators in region $A$, $Q_{A}=Q(x)=\int_{0}^{x}dx^{\prime}\rho(x^{\prime})=\rho_{0}x-\frac{1}{\pi}(\phi(x)-\phi(0))$ and $J_{A}=J(x)=\int_{0}^{x}dx^{\prime}j(x^{\prime})=\frac{uK}{\pi}(\theta(x)-\theta(0))=uK\tilde{J}_{A}=uK\tilde{J}(x)$.
Fractional charges and bipartite fluctuations.— The relevance of bipartite fluctuations measures can be understood from the ground-state correlation functions [2]
$\langle(\phi(x)-\phi(0))^{2}\rangle$ $=$ $K{\cal F}_{1}(x)$ (4)
$\langle(\theta(x)-\theta(0))^{2}\rangle$ $=$ $\frac{1}{K}{\cal F}_{1}(x)$
${\cal F}_{1}(x)$ $=$ $\frac{1}{2}\log\left(\frac{\alpha^{2}+x^{2}}{\alpha^{2}}\right),$
with $\alpha$ the lattice spacing or short-distance cutoff. These correlation functions will be precisely related to fluctuations of the charge $Q_{A}$ and current $\tilde{J}_{A}$. The function ${\cal F}_{1}$ measures the same information i.e. scaling properties as entanglement entropy at zero temperature [20]. Bipartitioning precisely means that the region $i\in[0;\frac{x}{\alpha}]$ is $A$ and the region $i\in[\frac{x}{\alpha};\frac{L}{\alpha}]$ is $B$. Bipartite charge fluctuations in a region of the wire with $x\gg\alpha$ (in the quantum limit $x\ll\beta u$ where $\beta=\frac{1}{k_{B}T}$) read [21]
${\cal F}_{Q(x)}$ $=$ $\langle Q(x)^{2}\rangle-\langle Q(x)\rangle^{2}=\frac{K}{\pi^{2}}\ln\frac{x}{\alpha}$
$=$ $\sum_{i,j\in[0;\frac{x}{\alpha}]}(\langle n_{i}n_{j}\rangle-\langle n_{i}\rangle\langle n_{j}\rangle),$
with $Q_{A}=Q(x)=\sum_{i=0}^{\frac{x}{\alpha}}c^{\dagger}_{i}c_{i}=\sum_{i=0}^{\frac{x}{\alpha}}n_{i}$. From the definitions above, we also have $n=n_{L}+n_{R}$ at position $i$ along the wire. In this formula, the measure is done at any time $t$ within the ground state. The symbol ${\cal F}_{Q}(x)$ can be thought of in a symmetric way i.e. ${\cal F}_{Q(x)}={\cal F}_{Q_{A}(x)}={\cal F}_{Q_{B}(L-x)}$ when the total charge $Q_{A}+Q_{B}$ is fixed. Bipartite charge fluctuations is related to the quantum Fisher information e.g. in the Kitaev p-wave superconductor [24] and Hubbard model [25]. The quantum Fisher information is measured in quantum spin chains through the dynamical spin susceptibility [26].
Here, we introduce bipartite fluctuations associated to the current $\tilde{J}(x)$ (which were not introduced before) from the fact that the correlator in $\theta$ also shows a logarithmic scaling:
${\cal F}_{\tilde{J}(x)}$ $=$ $\langle\tilde{J}(x)^{2}\rangle-\langle\tilde{J}(x)\rangle^{2}=\frac{1}{K\pi^{2}}\ln\frac{x}{\alpha}$ (6)
$=$ $\sum_{i,j\in A}(\langle\tilde{j}_{i}\tilde{j}_{j}\rangle-\langle\tilde{j}_{i}\rangle\langle\tilde{j}_{j}\rangle),$
with the corresponding current density on the lattice $j_{i}=it(c^{\dagger}_{i}c_{i-1}-c^{\dagger}_{i-1}c_{i})$ where $\tilde{j}_{i}=j_{i}/t$ for the 1D fermions, $\tilde{J}_{A}=\sum_{i\in A}\tilde{j}_{i}$ and $t$ the hopping strength (related to the hopping term); the Planck constant $\hbar=\frac{h}{2\pi}$ is set to unity. In the continuum limit, $\tilde{j}(x)=\frac{1}{K}(n_{R}(x)-n_{L}(x))$. Equivalently, ${\cal F}_{J(x)}=\langle J(x)^{2}\rangle-\langle J(x)\rangle^{2}=\frac{v_{F}^{2}}{K\pi^{2}}\ln\frac{x}{\alpha}$. The measure again is done at any time (usually, $t$ designates both the hopping amplitude and the time). Within our definitions, we can verify the identity $\frac{\partial}{\partial t}Q(x,t)=-\frac{\partial}{\partial x}J(x,t)$.
To reveal the chiral quasiparticles, we can then introduce the total charge in the region $A$ associated to left-moving and right-moving particles, $Q_{L}=\sum_{x^{\prime}\in A}n_{L}(x^{\prime})$ and $Q_{R}=\sum_{x^{\prime}\in A}n_{R}(x^{\prime})$, which are respectively related to $\phi_{L}$ and $\phi_{R}$ variables. Then, Eqs. (Bipartite Fluctuations and Charge Fractionalization in Quantum Wires) and (6) lead to
${\cal F}_{Q_{L}(x)}={\cal F}_{Q_{R}(x)}=\langle Q_{L}^{2}(x)\rangle-\langle Q_{L}(x)\rangle^{2}=\frac{K}{2\pi^{2}}\ln\frac{x}{\alpha}.$ (7)
This is also interpreted as $\frac{{\cal F}_{Q(x)}+{\cal F}_{K{\tilde{J}}(x)}}{4}={\cal F}_{\frac{Q(x)\pm K\tilde{J}}{2}}={\cal F}_{Q_{R}}(x)={\cal F}_{Q_{L}}(x)$ i.e. fluctuations associated to the chiral quasiparticles related to Eq. (3); this result assumes that $\langle\phi(x)\theta(0)\rangle\rightarrow cst.$ [2]. When adding Eqs. (Bipartite Fluctuations and Charge Fractionalization in Quantum Wires) and (6), then we obtain
$\pi^{2}\frac{({\cal F}_{Q(x)}+{\cal F}_{{\tilde{J}}(x)})}{2}=\frac{1}{K}(f_{R}^{2}+f_{L}^{2}){\cal F}_{1}(x).$ (8)
$\pi^{2}K^{2}\frac{({\cal F}_{Q(x)}+{\cal F}_{{\tilde{J}}(x)})}{2}=(f_{R}^{2}+f_{L}^{2})({\cal F}_{Q_{L}}(x)+{\cal F}_{Q_{R}}(x)).$ (9)
When we sum fluctuations of charge and current in the interacting electrons system this is equivalent to measure bipartite charge fluctuations associated to the chiral quasiparticles which carry a fractional charge and reveal $\frac{1}{K}(f_{L}^{2}+f_{R}^{2})=\frac{1}{2}(K+K^{-1})$ as a prefactor. The logarithmic function ${\cal F}_{1}$ can be seen as an entanglement measure of fractional charges. This can also be understood from the electron correlator in space $\langle c_{R}(x)c^{\dagger}_{R}(0)\rangle\propto\frac{1}{2\pi}\exp(-\frac{1}{K}(f_{R}^{2}+f_{L}^{2}){\cal F}_{1}(x))$, which can be verified from Eqs. (4) and standard definitions [2, 10]. It shows how an electron at $x=0$ is correlated to the position $x$ in space (at the same time) from the spread of the wavepackets at zero temperature. We also deduce
$\frac{\pi^{3}u}{4}\frac{\partial}{\partial x}({\cal F}_{Q}(x)+{\cal F}_{\tilde{J}}(x))=\frac{\pi u}{KL}(f_{R}^{2}+f_{L}^{2}).$ (10)
This is precisely the energy to add the pair of fractional charges along the quantum wire (with $x=\frac{L}{2}$). Here, we show the potential of Eq. (8) to measure bipartite fluctuations associated to mesoscopic interferences in a ring geometry. The signal detected from mesoscopic interferences can be written as the transmission probability [10] ${\cal T}=1-(T_{0}+T_{L})+\sqrt{T_{0}T_{L}}(e^{i\pi\varphi+iK_{1}-iK_{2}}+h.c.)$; $T_{0}$ and $T_{L}$ represent the transmission probabilities to enter the mesoscopic ring at $x=0$ and to escape from the ring at $x=L_{r}$, $\varphi=\Phi/\Phi_{0}$ is the magnetic flux with $\Phi_{0}$ the flux quantum. The phases $K_{1}$ and $K_{2}$ describe the dephasing from an electrical circuit with a resistance [13] and also the dephasing from interactions in the Luttinger liquid [10]. When we perform an ensemble of measures, this phase factor measures voltage fluctuations along one path $j$ of the ring through the identity $\langle e^{i{\cal K}_{j}}\rangle=e^{-\frac{\langle{\cal K}_{j}^{2}\rangle}{2}}$ [10]. The dephasing factor corresponds to voltage fluctuations integrated along a region of the ring [10]
$\langle{\cal K}_{j}^{2}\rangle=\frac{e^{2}}{\hbar^{2}}\int_{0}^{L_{r}}dx^{\prime}\int_{0}^{L_{r}}dx^{\prime\prime}\langle\delta V_{j}(x^{\prime})\delta V_{j}(x^{\prime\prime})\rangle.$ (11)
For Luttinger liquids, we can then rephrase the result at zero temperature in Ref. [10] as a correspondence with bipartite fluctuations in Eq. (9)
$\frac{1}{2}\langle{\cal K}_{j}^{2}(L_{r})\rangle=\pi^{2}\frac{({\cal F}_{Q(L_{r})}+{\cal F}_{{\tilde{J}}(L_{r})})}{2}-{\cal Y}$ (12)
where $\cal Y$} is the same quantity in the absence of interactions for $K=1$ i.e. $\pi^{2}\frac{({\cal F}_{Q(L_{r})}+{\cal F}_{{\tilde{J}}(L_{r})})}{2}-{\cal Y}=\left(\frac{1}{K}(f_{R}^{2}+f_{L}^{2})-1\right)\ln\frac{L_{r}}{\alpha}=2\gamma\ln\frac{L_{r}}{\alpha}$. Here, $2\gamma$ has the physical interpretation of a dimensionless resistance $r=\frac{Re^{2}}{h}$ and $x=L_{r}$ is the size of each path forming the ring. The form of fractional charges and the Luttinger parameter are measurable through the fit of this formula. The power-law form of mesoscopic interferences i.e. of $\langle{\cal T}\rangle$ is then a marker of the entangled nature of the two counter-propagating fractional wavepackets at zero temperature. When the right-moving charge (wavepacket) $f_{R}=\frac{1+K}{2}$ is measured at $x=L_{r}+\epsilon$ in the region $B$ then it is entangled with the left-moving fractional charge (wavepacket) which has moved backward and therefore is yet in the region $A$ inside the ring. The power-law decay of the interferences can be understood as the probability for the left-fractional charge $f_{L}=\frac{1-K}{2}$ to reach $x=L_{r}\sim ut$ at time $t$ i.e. $|\langle{\cal L}_{L}^{f_{L}}(x,t){\cal L}_{L}^{f_{L}}(0,0)\rangle|_{x=L_{r}\rightarrow ut}^{2}\propto\left(\frac{L_{r}}{a}\right)^{-2\gamma}$ [10].
Fractional Charges and Quantum Spin Chains.— Here, we introduce a correspondence with quantum spin chains from the Jordan-Wigner transformation at half-filling i.e. for $\langle n_{i}\rangle=\frac{1}{2}$. For spin polarized fermions with an interaction between nearest neighboring sites, the equivalent spin Hamiltonian is of the form of an XXZ chain
$H=-2t\sum_{j=1}^{N}(S_{j}^{x}S_{j+1}^{x}+S_{j}^{y}S_{j+1}^{y})+U\sum_{j=1}^{N}S_{j}^{z}S_{j+1}^{z},$ (13)
with the length of the wire $L=N\alpha$. We have the correspondences $S_{i}^{z}=c^{\dagger}_{i}c_{i}-\frac{1}{2}$, $S_{j}^{+}S_{j+1}^{-}=c^{\dagger}_{j}c_{j+1}$. The hopping term of fermions on the lattice becomes an $XY$ spin interaction (with equal couplings along $X$ and $Y$) and the interaction $U$ gives rise to an Ising interaction. When $U\ll 2t$, the low-energy theory of this quantum spin model at half-filling is therefore the Luttinger liquid in Eq. (1) [2, 28]. The electron charge density $n_{i}$ corresponds to the spin variable $S_{i}^{z}$, $n_{i}=S_{i}^{z}+\frac{1}{2}$ with $S_{i}^{z}=\pm\frac{1}{2}$ i.e. $\langle n_{i}\rangle=\frac{1}{2}$. The electron current density operator then precisely corresponds to
$j_{i}=t\tilde{j}_{i}=2t(S_{i}^{x}S_{i-1}^{y}-S_{i}^{y}S_{i-1}^{x})=2t{\cal O}_{i,i-1}^{xy-yx};$ (14)
the $\cal O$} operator coupling sites i,i1i,i-1 is the analogue of the $\tilde{j}_{i}=i(c^{\dagger}_{i}c_{i-1}-c^{\dagger}_{i-1}c_{i})$ current density introduced above. In the Letter [27], the authors introduce a representation of the current which can be written in terms of Majorana operators. We develop the correspondence with the ${\cal O}_{i,i-1}^{xy-yx}$ operator in the Supplementary Material [29].
For free fermions i.e. $U=0$, generalizing the derivations in Refs. [21], we find the exact formula from the $XY$ spin chain for bipartite current fluctuations [29]
${\cal F}_{\tilde{\cal J}(x)}$ $=$ $\sum_{i,j\in A}\langle{\cal O}_{i,i-1}^{xy-yx}{\cal O}_{j,j-1}^{xy-yx}\rangle-\langle{\cal O}_{i,i-1}^{xy-yx}\rangle\langle{\cal O}_{j,j-1}^{xy-yx}\rangle$ (15)
$=$ $\frac{1}{\pi^{2}}\left(\ln\frac{x}{\alpha}+\gamma+\ln 2+\frac{1}{2}\right).$
We remind the formula obtained for the charge fluctuations from the correlations of $S_{i}^{z}$ i.e. ${\cal F}_{Q(x)}=\frac{1}{\pi^{2}}(\ln\frac{x}{\alpha}+\gamma+\ln 2+1)$ [21, 29]. Adding interactions, this is identical to rescale the quantum fields as $\phi(x)\rightarrow\sqrt{K}\phi(x)$ and $\theta(x)\rightarrow\frac{1}{\sqrt{K}}\theta(x)$ in Eq. (1) compared to the free electron model such that we reproduce the leading form of the correlation functions in Eq. (6). From a correspondence with Bethe Ansatz, we have the identification for the Luttinger parameter $K=\frac{\pi}{2\arccos(-\frac{U}{2t})}$ [30].
In Figs. 1 and 2, we present DMRG results of the formula (9) with the quantum spin chain Hamiltonian at half-filling [31, 32]. For Periodic Boundary Conditions (PBC) charge fluctuations for a subsystem of size $x$ reveal a logarithmic behavior for $U/t=0$ up to and beyond $\sim 2$. We find the Luttinger parameter in Eq. (Bipartite Fluctuations and Charge Fractionalization in Quantum Wires) [21]. For Open Boundary Conditions (OBC) the charge fluctuations reveal an additional Friedel oscillating term [21]. We also verify the Luttinger parameter from the logarithmic term in the current fluctuations which is obtained from a precise protocol [29]; current fluctuations show an additional linear term which reveals the Mott transition in Fig. 2 inset. This method leads to the fit of fractional charges when adding the logarithmic terms of ${\cal F}_{\tilde{\cal J}}$ and ${\cal F}_{\cal Q}$ in Fig. 2. We also show a comparison with Bethe Ansatz.
Figure 1: Current fluctuations from DMRG with PBC for a wire with 300 sites, 20 sweeps. Data are in solid colored lines and the dashed lines represent a fit with a logarithmic term plus a linear term modulo a constant [29]. At weak $U$, results agree with the analytical form in Eq. (6) and when $U$ becomes substantial a linear form develops revealing the Mott transition.
Figure 2: Prefactor of the logarithmic term in $\pi^{2}({\cal F}_{Q}(x)+{\cal F}_{\tilde{J}}(x))$ with DMRG, same protocol as Fig. 1 and fit with the formula revealing the fractional charges in Eq. (8) with precisely $\frac{2}{K}(f_{R}^{2}+f_{L}^{2})=(\frac{1}{K}+K)$; the function $K$ in terms of $U/t$ is obtained from Bethe Ansatz [30]. (Inset) Increase of linear term at the Mott transition $(U=2t)$.
Mott physics.— In Ref. [33], we have shown that bipartite charge fluctuations are useful to study quantum phase transitions. We emphasize here that the bipartite current fluctuations are also a powerful and precise tool to detect the vicinity of insulating or Mott phases associated to the decay of the logarithmic term of ${\cal F}_{\tilde{\cal J}}$ in Fig. 2. In the spin chain correspondence, this occurs when $U>2t$ or $K<\frac{1}{2}$ such that the interaction term proportional to $(c^{\dagger}_{R}c_{L})^{2}+h.c.$ in the quantum field theory, i.e. the Sine-Gordon term $\cos(4\phi\sqrt{K})$ coming from $S_{j}^{z}S_{j+1}^{z}$ becomes relevant [28], i.e. the mode $\phi(x)$ becomes fixed to a uniform value associated to this insulator i.e. a classical antiferromagnet in the spin language. The Heisenberg spin chain corresponds to $U=2t$ i.e. $K=\frac{1}{2}$. Fluctuations in the charge sector remain very small for $U>2t$ [29]. The electron Green’s function decreases exponentially in space $\sim\exp-\frac{x}{\xi}$ as a result of the charge density wave formation and the linear term in the bipartite current fluctuations $\frac{x}{\xi}$ i.e. give an access to the correlation length $\xi$ or inversely the energy gap. A fit with inset of Fig. 2 gives the estimate $\frac{\alpha}{\xi}\sim e^{\frac{U}{t}1.4533}$.
Localization effect from the interface and Jackiw-Rebbi topological model.— The Jackiw-Rebbi model [34] is a general model of interfaces with two media showing a potential difference such that it admits a zero-energy bound state in the solution of the quantum field theory [35]. The zero-energy bound state occurs even if the potential is smoothly varying around the interface located at $x=\frac{L}{2}$ i.e. with a potential of the form $\sim\Delta\tanh\left(\frac{x-\frac{L}{2}}{\alpha}\right)$ [34, 36]. Here, we ask two questions: can such a bound state be observable with our tool e.g. bipartite current fluctuations? Can the zero-energy solution coexist with the fractional charges of the Luttinger liquid? We show below that our analysis can lead to positive answers. Our tool can then be useful to locate interfaces as well.
We introduce the variable $y=x-\frac{L}{2}$ such that the bound state will be localized around $y=0$. Looking for eigenstates resolved in energy $\epsilon$, it is useful to introduce
$c_{\epsilon}^{\dagger}=\int_{-\frac{L}{2}}^{\frac{L}{2}}dy\chi_{\epsilon}(y)c^{\dagger}(y),$ (16)
associated to the Hamiltonian $\sum_{\epsilon}\epsilon c^{\dagger}_{\epsilon}c_{\epsilon}$. It allows us to show how the eigenstate equation $H,c^{\dagger}_{\epsilon}]=\epsilon c^{\dagger}_{\epsilon}$ turns into [29]
$\epsilon\chi_{\epsilon}(y)=-2iv_{F}\partial_{y}\chi_{\epsilon}(y)-i2\Delta\tanh\left(\frac{y}{\alpha}\right)\chi_{\epsilon}(-y).$ (17)
We derive a proof of such an equation formulating a correspondence towards a quantum field theory of a 1D superconducting wire with s-wave (or d-wave) symmetry or a Luther-Emery model [37, 38, 39]. Eq. (17) admits the zero-energy state [29]
$\chi_{0}(y)=\left(\frac{1}{\sqrt{\pi}}\sqrt{\frac{\Delta}{2t}}\right)^{\frac{1}{2}}e^{-\frac{\Delta}{2t}\ln\cosh\frac{|y|}{\alpha}}.$ (18)
We assume half-filling such that $v_{F}=2t$. First, we verify the occurrence of the bound state in the density probability and in the bipartite current fluctuations at weak $U$; see Fig. 3. In Fig. 3 with OBC, we clearly observe a logarithmic term at small $x$ for current fluctuations, traducing the Luttinger liquid, coexisting with the bound state response at $x=\frac{L}{2}$. The prefactor of the logarithmic term decays as $\sim(1-\left(\frac{\Delta}{2t}\right)^{2})$ at small to intermediate $U$, which corresponds to the probability for an electron to cross the interface or to the fractional charges to spread on the whole sample. This result shows that fractional charges can yet develop a Luttinger liquid in the presence of the interface. For OBC charge fluctuations show pronounced oscillations as a function of distance.
From bosonization, the interface produces a backscattering term $\sim\frac{\Delta}{2\pi\alpha}\tanh\left(\frac{x-\frac{L}{2}}{\alpha}\right)\cos(2\phi)$ in the Hamiltonian density (close to the interface) [2]. It goes to zero precisely when $x=\frac{L}{2}$ and for $x\neq\frac{L}{2}$ it takes a dual form as the p-wave pairing term [40] in the Kitaev p-wave superconducting wire [41], if we flip the role of the charge field $\phi$ and current field $\theta$. From this sense, this bound state can be seen as topological [35], i.e. is stable in the presence of interactions. We emphasize here that the peaks in the density probability and in the current fluctuations remain visible for strong interactions $U\sim t$ and similarly for the logarithmic scaling at small $x$ [29].
Figure 3: Bipartite Current Fluctuations for OBC at $U=0.2t$ with 300300 sites, for various values of $\Delta$. (Inset) Density probability (not normalized) at small $U$ e.g. at $U=0$.
Conclusion.— We have derived the relation between fractional charges in 1D quantum wires and bipartite chiral charge fluctuations, asumming bipartite charge and current fluctuations. This also provides a way to measure them through the mesoscopic interferences at zero temperature. We have shown a spin chain correspondence and verified the occurrence of fractional charges through DMRG. We have also shown how bipartite current fluctuations represent a useful marker to localize Mott transitions through a linear term. Introducing a smoothly varying potential at the interface between the two sides of the wire, it is yet possible to detect a localized bound state in the presence of interactions coexisting with the fractional charges. Bipartite charge, spin fluctuations and also bipartite current fluctuations may be measured from correlation functions with a double summation on space. We observe progress in developing tools e.g. in ultra-cold atoms [42]. Similarly, coupling a metallic gate along a region of the wire would also allow to resolve information on charge dynamics similarly as in a quantum dot [43]. Our bipartite fluctuations [21] may also be measured from the dynamical spin correlations similarly as for the measure of the quantum Fisher information [25, 26]. Tools are also developed to measure fractional charges in fractional quantum Hall states [14, 15, 16, 17, 22, 23].
Magali Korolev is grateful to Ecole Polytechnique for the kind support of her PhD thesis. The numerical results are obtained with the Cluster Cholesky at Ecole Polytechnique.